OpenStudy (anonymous):

The angles shown below are supplementary: A pair of supplementary angles is shown. One angle measures 10x and the other angle measures 60 degrees. What is the value of x? 6° 9° 12°

OpenStudy (anonymous):

@amistre64 last one i promise sorry (:

OpenStudy (amistre64):

well we need a definition, what does it mean to be complementary angles?

OpenStudy (amistre64):

ugh, supplementary angles that is

OpenStudy (amistre64):

might as well find the definition of both, knowing what complementary and supplementary angles are wont hurt :)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

two angles are supplimentary when they add up to 180 degrees

OpenStudy (amistre64):

thats a good definition, so all we need to do is add the angles and equate it to 180 10x + 60 = 180 .... what should x be?

OpenStudy (anonymous):

arent we supposed to subtract something

OpenStudy (amistre64):

sure, it looks like that +60 is getting in the way, lets subtract it ... from BOTH sides 10x + 60 = 180 - 60 -60 -------------- 10x + 0 = 120 so, now we can focus on: 10x = 120 what would you attempt now?

OpenStudy (amistre64):

10, times what, is 120?

OpenStudy (anonymous):

12

OpenStudy (amistre64):

the 10x = 120, when x=12

OpenStudy (amistre64):

and 10x + 60 = 180, when x=12

OpenStudy (anonymous):

so 12 degrees is the answer (:

OpenStudy (amistre64):

well, not 12 degrees, 10x degrees, x is just a number that makes that value of 10x degrees equal ,120 degrees

OpenStudy (amistre64):

x=12 will suffice

OpenStudy (anonymous):

so then what is the answer ?

OpenStudy (amistre64):

hmm, i just noticed that they choose to put it as 12 degrees. old tired eyes .... sure, lets pick the one thats a 12, if you want to tack on a degree notation, its fine apparently :)

OpenStudy (anonymous):

ok Thank you so much ! (:

OpenStudy (amistre64):

yep